OpenStudy (anonymous):

a. Solve 4x^2=20 b. Find the x-intercepts of f(x) =4x^2-20 a. What are the solutions? x=_____. b. What are the intercepts? The x-intercepts are_____.

OpenStudy (anonymous):

4x^2=20 4x^2-20 =0 divide by4 x^2-5=0 x=-+sqrt5

OpenStudy (anonymous):

when f(x)=0, x1=sqrt5 and x2=-sqrt5 x intercepts when x=0, fx()=-20 y intercept

OpenStudy (anonymous):

oops sorry when x=0, f(x)=-20 y intercept